# Need Help? Designing Proactive AI Assistants for Programming

### v1

@@ Abstract
While current chat-based AI assistants primarily operate reactively, responding only when prompted by users, there is significant potential for these systems to proactively assist in tasks without explicit invocation, enabling a mixed-initiative interaction. This work explores the design and implementation of proactive AI assistants powered by large language models. We first outline the key design considerations for building effective proactive assistants. As a case study, we propose a proactive chat-based programming assistant that automatically provides suggestions and facilitates their integration into the programmer’s code. The programming context provides a shared workspace enabling the assistant to offer more relevant suggestions. We conducted a randomized experimental study examining the impact of various design elements of the proactive assistant on programmer productivity and user experience. Our findings reveal significant benefits of incorporating proactive chat assistants into coding environments and uncover important nuances that influence their usage and effectiveness.
@@
Figure 1: The implementation of a proactive chat assistant for programming. We introduce a proactive assistant that takes in the programmer’s context, which includes the current code and user message history, generates a set of suggestions, which include a summary, implementation, and explanation of implementation, and then determines whether it is timely to show the user in the interactive interface.
1 Introduction
@@
Chat-based AI assistants such as ChatGPT [1] or Claude [2], enable people to accomplish some of their tasks via prompting: the user crafts a message with sufficient context about their task and then receives an AI response, they continue iterating this process until they are satisfied with the result. Common use cases of chat-based AI assistance include writing tasks [3, 4], idea generation [5], programming tasks [6, 7], and web navigation among other tasks [8, 9]. Interacting with these assistants requires two points of effort from the human: first, porting over the workspace context, e.g., copying the document or code file, and second, describing the task in natural language. Assuming the assistant had access to the user’s working context, a natural question is whether it could infer how it can help the user with the task they are trying to solve. If the assistant can reliably infer the user’s intent, generating a response to the task and suggesting it automatically might be advantageous. The assistant would now be proactive and result in a mixed-initiative interaction [10]. In this paper, we attempt to design effective proactive chat-based assistants powered by large language models (LLMs) for programming.
@@
Virtual assistants have been explored in many commercial products including Apple Siri and Amazon Alexa [11]; in fact, one of the most well-known and first proactive virtual assistants is Microsoft’s Office Assistant, with the English version commonly referred to as Clippy. The Office Assistant would proactively offer assistance based on inference of the user’s intent from their Word document. While the user experience with Clippy was less than ideal [12], more recent examples of proactive assistants in the form of autocompletion for text-based tasks have been more successful. For instance, GitHub Copilot [13] provides in-line code suggestions with LLMs. Copilot has been well received by programmers and randomized studies have shown increased programmer productivity [14, 15]. However, autocomplete-based proactive assistance only considers local suggestions based on the user’s cursor position. In contrast, a chat-based proactive assistant could suggest modifications to or point out bugs in earlier parts of the code, in addition to being able to complete the code. Moreover, the particular design details of a proactive assistant can significantly impact its effectiveness.
@@
We begin by outlining design considerations for designing chat-based proactive assistants with LLMs to address challenges of proactivity [16] by building on previous guidelines for human-AI interaction and mixed-initiative interaction [10, 17, 4]. Key considerations include supporting efficient evaluation and utilization of proactive suggestions and timing suggestions based on context. Given the success of LLM-based assistants for programming, the programming domain presents a good opportunity to study proactivity. Since the context of the domain is well-defined—consisting of the code file(s) in a programmer’s integrated development environment (IDE), integrating the proactive assistant into the programmer’s IDE enables us to overcome the challenge of observing the human workspace. Based on these design considerations, we build a proactive assistant for programming and integrate it into the chat interface within a code editor, as shown in Figure 1. The chat interface operates like a standard interface where users can type messages and receive responses; however, the assistant can now send proactive messages without a corresponding user request. The timing of the proactive suggestions depends on the human’s recent activities in the editor and their interactions with the assistant responses. Proactive messages are distinguished from regular assistant messages, with the former only displaying a high-level description unless the user chooses to expand the message to see a suggested implementation or explanation. Finally, each assistant message can be integrated into the code via a “preview” button highlighting the changes the message will make.
@@
We conduct a user study to evaluate programmer usage of our proactive assistant against a baseline ChatGPT-like assistant. For our experiments, we use the web-based code editor RealHumanEval [18] as it allows us to easily conduct a controlled study to compare baseline and proactive conditions. To understand the effects of changing different design decisions, we investigate multiple variants of our proactive assistants, which differ in the timing of the suggestions and the ability to integrate suggestions into the programmer’s code. Our study demonstrates the benefits of proactivity in chat assistants for code as proactive assistants increase the number of tasks completed by an average of 12-18%. However, our study also shows that proactive assistants should be designed carefully as small changes can lead to differing user experiences. For example, 80-90% participants preferred the proactive assistant over the baseline chat assistant but only half the participants preferred proactive assistants that provided more frequent suggestions, despite productivity gains. We examined the telemetry and found that providing too many suggestions can saturate the utility of suggestions to users, leading them to copy suggested code into their editor less often. Furthermore, we observed that people tend to rely on proactive assistant capabilities when available, including the ability to incorporate code manually, rather than their manual efforts. Our findings demonstrate the potential of proactive assistants to improve programmer productivity and experience and highlight important design decisions that impact these downstream metrics.
@@
  • A controlled user study characterizing the benefits of our proactive chat assistant on downstream user metrics—e.g., productivity and user experience (Section 6.1 and Section 6.2), highlighting the potential for proactivity to be included in future coding assistants, as well as demonstrating the potential negative effects of changing certain proactive features. We also demonstrate how changes in proactive assistant capabilities lead to different interaction patterns. (Section 6.3).
2 Related Work
2.1 The Ecosystem of Programming Assistants
@@
A growing number of tools powered by large language models (LLMs)—i.e., programming assistants—are available to developers to generate or edit code and answer queries. Programmers are increasingly writing code with AI assistants like Github Copilot [13] and Cursor [19] and are using chat assistants like ChatGPT [1] or Claude [2] in place of online Q&A communities like Stack Overflow [6]. Programming assistants typically involve one of two types of support: autocomplete suggestions are used to quickly write more code based on the programmer’s current code context, while chat dialogue can help answer questions that range from debugging errors to explaining documentation. Programming assistants surrounding code completion have been the focus of prior studies, which provide insight into how people use LLM-generated code completions [14, 7, 20, 21, 15, 22] and how to improve completions by providing explanations [23] and determining when to best show [24, 25]. While there has been less focus on investigating the use of chat assistants for programming, it is important to note that they complement assistance provided by code completions [26] and are the focus of this work.
@@
Prior studies investigating chat dialogue for programming assistance have typically considered the setting where people initiate questions to the assistant [27, 28, 29, 6, 30, 31]. In particular, Chopra et al. [28] found significant time is spent constructing prompts and expressing their context to ChatGPT. As such, people did not enjoy being “slowed down” through this process. Additionally, Nam et al. [30] noted that people often had difficulty finding a good prompt for the chat assistant to provide the desired response. Further, Mozannar et al. [18] highlighted how the burden remains for programmers to incorporate an assistant response into their code. These observations highlight the hurdles to a streamlined integration of chat assistants like ChatGPT in a developer’s day-to-day workflow: the need to provide sufficient context about the developer’s problem and for developers to manually input and decide what they want to ask and then act upon the assistant response. Even while some programming assistants (e.g., Cursor [19], Github Copilot [13]) have begun building chat assistants into the integrated development environment (IDE), they are still reactive. In this work, we study whether incorporating proactivity into chat assistants can improve productivity and user experience. Our work aligns with existing tools for programming that provide chat support in that we still allow developers to ask questions when they so choose. Different from prior work, our proactive chat assistant periodically recommends suggestions and allows developers to integrate those suggestions into their code.
2.2 Background on Proactive Assistants
@@
Research on proactive assistants has appeared in many forms, ranging from physical robots [32, 33, 34] to virtual chat assistants on phones or computer applications [35, 36, 37, 38]. Given our target application of programming, we focus our discussion on virtual chat assistants. Even within proactive chat assistants, the goals of prior proactive assistants vary depending on the context they are used. For example, support-based assistants use proactivity to provide motivation and continued dialogue [36, 38], while education-related assistants can proactively offer relevant support or explanations for students [39]. Most relevant to our domain are the set of information-finding assistants that leverage proactive messaging to provide additional and useful information [40, 41] and planning assistants can proactively help people reason about their decisions and identify overlooked alternative decisions or rationales [37]. The goal of our proactive assistant is to improve a programmer’s productivity while maintaining a good user experience; it requires blending multiple goals of existing proactive assistants. In Section 3, we discuss further how these downstream goals inform the design considerations for our proactive assistant.
@@
Many prior deployed proactive assistants have been faced with significant negative reaction because the actual capabilities of these systems do not meet user expectations [11, 42], as many of the systems relied on pre-set messaging or simple models [35, 36]. With the advent of modern LLMs, and their growing usage as the backbone of agents [43, 44, 45], there is significant potential to revisit how we design more capable proactive assistants. In recent work, ComPeer [38] leveraged GPT-4 to build a proactive assistant for mental health support that handles and acts on a longer term memory of user dialogue. Aside from mental health applications, we believe LLMs can be particularly well suited for proactive assistants for productivity goals, particularly when they are already being used as regular chat assistants (e.g., ChatGPT). The increased model capability and ability to handle long contexts would be suitable for capturing complex environments in which they are completing tasks, which has been relatively unexplored in prior chat assistant literature relating to productivity [16]. In Section 4, we discuss the design of our proactive system which leverages off-the-shelf state-of-the-art LLMs to propose suggestions based on the current code context and even incorporate these suggestions into user code.
3 Design Considerations for Proactive Coding Assistants
@@
To formulate design considerations for the core functionality and behavior of the proactive assistant, we revisit existing guidelines on designing mixed-initiative user interfaces and human-AI interactions [10, 17] as well as a prominent survey of known challenges of designing proactive chat assistants [16]. We highlight five design considerations that merit careful consideration to capture the benefits and challenges of proactivity.
@@
Harnessing the benefits of proactivity. Horvitz [10] emphasizes the importance of “developing significant value-added automation” when developing effective mixed-initiative systems. Proactivity can address the limitations faced by programmers when interacting with existing, reactive chat assistants for coding contexts. As discussed in Section 2.1, the value-add of introducing proactive suggestions would be an improvement in programmer productivity (e.g., in terms of time and effort) and in the user experience (e.g., reducing perceived “slow-down”), as compared to a non-proactive chat assistant [28, 30, 18]. To ensure that programmers can effectively use proactive suggestions, we propose the following two design considerations:
@@
(1) Support efficient evaluation. Proactive suggestions should be provided in a way such that programmers can efficiently interact with them. This involves providing a manageable amount of information to the user about the suggestions so that they can quickly decide if they are worthwhile. This might involve hierarchically scaffolding the suggestions so the programmer can obtain more information about the suggestion if needed.
@@
(2) Support efficient utilization. If the user decides the suggestion is useful, the proactive assistant should make it easy for the programmer to utilize. For example, the proactive assistant should make it as easy as possible for them to take action corresponding to the suggestion and incorporate it into the programmer’s code. If the programmer decides the suggestion is not useful, they should be able to dismiss the assistant easily.
@@
Handling challenges of proactivity. Prior works on proactive chat assistants have demonstrated that there is considerable nuance in developing a proactive chat assistant [16], as irrelevant and untimely proactive messages may compromise the success of the interaction timing [46, 35, 47, 48]. We incorporate these challenges in our design considerations:
@@
(3) Show contextually relevant suggestions. Proactive chat assistants are increasingly built into complex environments where users often perform multiple different tasks. For example, programming tasks can range from implementation to debugging and testing, to name a few, that require different interventions and information from the assistant. Proactive assistants should make sure to surface information that is relevant to the current context and recent user interactions and messages.
(4) Time suggestions based on context. Programmers perform various tasks during their work and are often in a state of flow [49]. The proactive assistant should consider what the programmer is currently doing and their workflow before it provides a suggestion [20]. For example, showing suggestions while the user is actively coding can disrupt the programmer’s flow and lead to disruptions. Horvitz [10] discusses the importance of fallback behavior which can allow for a reduced frequency of suggestions.
@@
(5) Incorporating user feedback. The proactive assistant should adapt its suggestions and timing as it interacts with the user. The user should be able to accept and reject proactive suggestions, and the decision to accept or reject suggestions should influence future decisions of the proactive assistant. This is especially important as different programmers may want different levels of proactive help from the assistant.
@@
In the following section, we introduce an implementation of a proactive assistant for programming building on the design considerations above.
4 Proactive Assistant Implementation

### v2

@@ Abstract.
While current chat-based AI assistants primarily operate reactively, responding only when prompted by users, there is significant potential for these systems to proactively assist in tasks without explicit invocation, enabling a mixed-initiative interaction. This work explores the design and implementation of proactive AI assistants powered by large language models. We first outline the key design considerations for building effective proactive assistants. As a case study, we propose a proactive chat-based programming assistant that automatically provides suggestions and facilitates their integration into the programmer’s code. The programming context provides a shared workspace enabling the assistant to offer more relevant suggestions. We conducted a randomized experimental study examining the impact of various design elements of the proactive assistant on programmer productivity and user experience. Our findings reveal significant benefits of incorporating proactive chat assistants into coding environments, while also uncovering important nuances that influence their usage and effectiveness.
AI-assisted Programming, Proactivity, Mixed-Initiative Interaction
journalyear: 2025copyright: ccconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714002isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing User studiesccs: Software and its engineering Collaboration in software development
@@
Figure 1. The implementation of a proactive chat assistant for programming. We introduce a proactive assistant that takes in the programmer’s context, which includes the current code, user message history, and optionally terminal output, generates a set of suggestions, which include a summary, implementation, and explanation of implementation, and then determines whether it is timely to show the user in the interactive interface.
[…]
The implementation of a proactive chat assistant for programming. The proactive assistant takes in the programmer’s context, which includes the current code, user message history, and optionally terminal output, generates a set of suggestions, which include a summary, implementation, and explanation of implementation, and then determines whether it is timely to show the user in the interactive interface.
1. Introduction
@@
Chat-based AI assistants such as ChatGPT (OpenAI, 2022) or Claude (Anthropic, 2023), enable people to accomplish some of their tasks via prompting: the user crafts a message with sufficient context about their task and then receives an AI response, they continue iterating this process until they’re satisfied with the result. Common use cases of chat-based AI assistance include writing tasks (Lee et al., 2022, 2024), idea generation (Shaer et al., 2024), programming tasks (Xiao et al., 2023; Mozannar et al., 2024a), and web navigation among other tasks (Lin et al., 2022; Yao et al., 2024). Interacting with these assistants requires two points of effort from the human: first porting over the workspace context, e.g., copying the document or code file, and second describing the task in natural language. Assuming the assistant had access to the working context of the user, a natural question is whether it could infer the task they want to solve. If the assistant can reliably provide suggestions relevant to the user’s task, then it might be advantageous for it to generate suggestions automatically. The assistant would now be proactive and result in a mixed-initiative interaction (Horvitz, 1999). In this paper, we attempt to design effective proactive chat-based assistants powered by large language models (LLMs) for programming.
@@
Virtual assistants have been explored in many commercial products including Apple Siri and Amazon Alexa (Luger and Sellen, 2016); in fact, one of the most well-known and first proactive virtual assistants is Microsoft’s Office Assistant, with the English version commonly referred to as Clippy. The Office Assistant would proactively offer its assistance based on inference of the user actions in their Word document. While the user experience with Clippy was less than ideal (Baym et al., 2019), more recent examples of proactive assistants in the form of autocompletion for text-based tasks have been more successful. For instance, GitHub Copilot (Github, 2022) provides in-line code suggestions with LLMs. Copilot has been well received by programmers and randomized studies have shown increases in programmer productivity (Vaithilingam et al., 2022; Peng et al., 2023). However, autocomplete-based proactive assistance only considers local suggestions based on the user’s cursor position. In contrast, a chat-based proactive assistant could suggest modifying the entire code or pointing out bugs in earlier parts of the code, in addition to being able to complete the code. Moreover, it is clear that the particular design details of a proactive assistant are crucial to its effectiveness.
@@
We begin by outlining design considerations for designing chat-based proactive assistants with LLMs to address challenges of proactivity (Chaves and Gerosa, 2021) by building on previous guidelines for human-AI interaction and mixed-initiative interaction (Horvitz, 1999; Amershi et al., 2019; Lee et al., 2024). Key considerations include supporting efficient evaluation and utilization of proactive suggestions as well as timing of suggestions based on context. The domain of programming presents a good opportunity to study proactivity given the success of LLM-based assistants for programming and a well-defined work context, which consists of the code file(s). This enables us to overcome the challenge of observing the human’s workspace by integrating the assistant into the programmer’s integrated development environment (IDE). Based on these design considerations, we build a proactive assistant for programming and integrate it into a chat interface inside a code editor, as shown in Figure 1. The chat interface operates like a standard interface where the user can type messages and receive responses; however, the assistant can now send proactive messages without a corresponding user request. The timing of the proactive suggestions depends on the human’s recent activities in the editor and their interactions with the assistant responses. The proactive messages are distinguished from regular assistant messages, with the former only displaying a high-level description of the message unless the user chooses to expand the message. Finally, each assistant message can be integrated into the code via a “preview” button that highlights the changes the message will make to the code.
@@
We perform a controlled user study where we evaluate programmer usage of our proactive assistant against a baseline ChatGPT-like assistant. We use the web-based code editor RealHumanEval (Mozannar et al., 2024c) to experiment as it allows us to easily conduct studies. In the study, we include multiple conditions varying different aspects of the proactive assistant, including the timing of the suggestions and the ability to integrate suggestions into the programmer’s code, to understand the effects of changing different design decisions. Our study demonstrates the benefits of proactivity in chat assistants for code as proactive assistants increase the number of tasks completed by 12-18%. However, our study also shows that proactive assistants need to be designed carefully as small changes can lead to differing user experiences—for example, increasing the frequency of suggestions can negatively impact user experience, reducing participant preference for the proactive assistant over baseline by half, despite productivity gains. Additionally, small changes in proactive assistant capabilities, including the ability to integrate suggestions, can significantly change user interaction and coding patterns.
@@
  • A controlled user study characterizing the benefits of our proactive chat assistant on downstream user metrics (e.g., productivity and user experience), highlighting the potential for proactivity to be included in future coding assistants, as well as demonstrating the potential negative effects of changing certain proactive features (Section 6.1 and Section 6.2). We also analyze fine-grain user behavior when interacting with different proactive assistants (Section 6.3).
2. Related Work
2.1. The Ecosystem of Programming Assistants
@@
A growing number of tools powered by large language models (LLMs)—i.e., programming assistants—are available to developers to generate or edit code and answer queries. Programmers are increasingly writing code with AI assistants like Github Copilot (Github, 2022) and Cursor (Cursor, 2023) and are using chat assistants like ChatGPT (OpenAI, 2022) or Claude (Anthropic, 2023) in place of online Q&A communities like Stack Overflow (Xiao et al., 2023). Programming assistants typically involve one of two types of support: autocomplete suggestions are used to quickly write more code based on the programmer’s current code context, while chat dialogue can help answer questions that range from debugging errors to explaining documentation. Programming assistants surrounding code completion have been the focus of prior work, providing insight into how people use LLM-generated code completions (Vaithilingam et al., 2022; Mozannar et al., 2024a; Barke et al., 2023; Prather et al., 2023; Peng et al., 2023; Cui et al., 2024) and improving completions by providing explanations (Yan et al., 2024) and determining when to best show (Vasconcelos et al., 2023; Mozannar et al., 2024b). While there has been less focus on investigating the use of chat assistants for programming, it is important to note that they complement assistance provided by code completions (Liang et al., 2023) and are the focus of this work.
@@
Prior studies investigating chat dialogue for programming assistance have always looked at the setting where people initiate questions to the assistant (Ross et al., 2023; Chopra et al., 2023; Kazemitabaar et al., 2023; Xiao et al., 2023; Nam et al., 2024; Kazemitabaar et al., 2024). In particular, Chopra et al. (2023) found that people spent a significant amount of time constructing prompts and gathering and expressing their context to ChatGPT and did not enjoy being “slowed-down” through this process. Additionally, Nam et al. (2024) noted that people often had a hard time finding a good prompt for the chat assistant that could give them the desired response. Further, Mozannar et al. (2024c) highlighted how the burden remains for programmers to appropriately incorporate an assistant response into their code. These observations highlight the hurdles to a streamlined integration of chat assistants like ChatGPT in a developer’s day-to-day workflow: the need to provide sufficient context about the developer’s problem and the need for developers to manually input and decide what they want to ask and then act upon the assistant response. Even while some programming assistants (e.g., Cursor (Cursor, 2023), Github Copilot (Github, 2022)) have begun building chat assistants into the integrated development environment (IDE), they still require human input. In this work, we evaluate whether incorporating an aspect of proactivity into chat assistants can lead to more benefits in terms of productivity and user experience. Our work aligns with existing tools for programming that provide chat support in that we still allow developers the option to ask questions when they so choose. Different from prior work, our proactive chat assistant also periodically recommends suggestions and even allows developers to integrate those suggestions into their code.
While the focus of this work is on designing proactive chat-based assistants, we overview other forms of proactive assistance in the existing programming literature. As mentioned in Section 1 and discussed extensively in Section 2.1, the most well-known and widely used form of proactive assistance is in-line code completions (e.g., GitHub Copilot (Github, 2022)(Liang et al., 2023). Aside from proactive code completions, multiple related works have studied proactive assistance to address various issues including fixing error messages and understanding LLM generations. For example, a set of more traditional approaches explored the use of adaptive feedback—drawn from existing examples or code repair tools—to help students resolve error messages in submitted programs (Hartmann et al., 2010; Head et al., 2017; Ahmed et al., 2020). Recent approaches have explored live programming techniques to “preview” the outcomes of AI-generated code using projection boxes (Lerner, 2020; Ferdowsi et al., 2024). Our work investigates whether LLM-powered proactive assistants can address multiple issues at once. We return to these studies and other aforementioned studies on chat-based assistance for programming to discuss how our findings inform the design of proactive programming assistants in Section 7.1.3.
2.2. Background on Proactive Assistants
@@
Research on proactive assistants has appeared in many forms, ranging from physical robots (Zhang et al., 2015; Baraglia et al., 2016; Peng et al., 2019) to virtual chat assistants on phones or computer applications (Liao et al., 2016; Fitzpatrick et al., 2017; Gu et al., 2024; Liu et al., 2024b). Given our target application of programming, we focus our discussion on chat assistants. Even within proactive chat assistants, the goals of prior proactive assistants vary depending on the context they are used. For example, support-based assistants use proactivity to provide motivation and continued dialogue (Fitzpatrick et al., 2017; Liu et al., 2024b), while education-related assistants can proactively offer relevant support or explanations for students (Winkler and Roos, 2019). Most relevant to our domain are the set of information-finding assistants that leverage proactive messaging to provide additional and useful information (Avula et al., 2018; Hu et al., 2024) and planning assistants can proactively help people reason about their decisions and identify overlooked alternative decisions or rationales (Gu et al., 2024). The goal of our proactive assistant is to improve a programmer’s productivity while maintaining a good user experience; it requires blending multiple goals of existing proactive assistants. In Section 3, we discuss further how these downstream metrics inform the design considerations for our proactive assistant.
@@
Many prior deployed proactive assistants have failed or received significant negative reaction because the actual capabilities of these systems do not meet user expectations (Luger and Sellen, 2016; Meurisch et al., 2020), as many of the systems relied on pre-set messaging or simple models (Liao et al., 2016; Fitzpatrick et al., 2017). With the advent of modern LLMs, and their growing usage as the backbone of agents (Zhou et al., 2023; Park et al., 2023; Yang et al., 2024), there is significant potential to revisit how we design more capable proactive assistants. In recent work, ComPeer (Liu et al., 2024b) leveraged GPT-4 to build a proactive assistant for mental health support that handles and acts on a longer term memory of user dialogue. Aside from mental health applications, we believe LLMs can be particularly well suited for proactive assistants for productivity goals, particularly when they are already being used as regular chat assistants (e.g., ChatGPT). The increased model capability and ability to handle long contexts would be suitable for capturing complex environments in which they are completing tasks, which has been relatively unexplored in prior chat assistant literature relating to productivity (Chaves and Gerosa, 2021). In Section 4, we discuss the design of our proactive system which leverages off-the-shelf state-of-the-art LLMs to propose suggestions based on the user environment and even incorporate these suggestions into user code.
3. Design Considerations for Proactive Coding Assistants
@@
To formulate design considerations for the core functionality and behavior of the proactive assistant, we revisit existing guidelines on designing mixed-initiative user interfaces and human-AI interactions (Horvitz, 1999; Amershi et al., 2019) as well as a survey of known challenges of designing proactive chat assistants (Chaves and Gerosa, 2021). We highlight five design considerations that merit further attention to capture the benefits and challenges of proactivity.
@@
Harnessing the benefits of proactivity. Horvitz (1999) emphasizes the importance of “developing significant value-added automation” when developing effective mixed-initiative systems. Given prior work highlighting the limitations of existing chat assistants for coding contexts (Chopra et al., 2023; Nam et al., 2024; Mozannar et al., 2024c) (as discussed in Section 2.1), the value-add of introducing proactivity would be an improvement in programmer productivity (e.g., in terms of time and effort) and in the user experience (e.g., reducing perceived “slow-down”), as compared to a non-proactive chat assistant. Concretely, we propose the following two design considerations:
@@
(1) Support efficient evaluation. Proactive suggestions should be provided in a way such that programmers can efficiently interact with them. This involves providing a manageable amount of information to the user about the suggestions so that they can quickly decide if they are worthwhile. This might involve hierarchically scaffolding the suggestions so that the programmer can obtain more information about the suggestion if needed.
@@
(2) Support efficient utilization. If the user decides the suggestion is useful, the proactive assistant should make it easy for the programmer to utilize it. For example, the proactive assistant should make it as easy as possible for them to take action corresponding to the suggestion and incorporate it into the programmer’s code. If the programmer decides the suggestion is not useful, they should be able to easily dismiss the assistant.
@@
Handling challenges of proactivity. Prior works on proactive chat assistants have demonstrated that there is considerable nuance in developing a proactive chat assistant (Chaves and Gerosa, 2021), as untimely and irrelevant proactive messages may compromise the success of the interaction timing (Silvervarg and Jönsson, 2013; Liao et al., 2016; Toxtli et al., 2018; Tallyn et al., 2018). We incorporate these challenges in our design considerations:
@@
(3) Show contextually relevant suggestions. Proactive chat assistants are increasingly built into complex environments where users often perform a number of different tasks. For example, programming tasks can range from implementation to debugging and testing, just to name a few, which all would require different interventions and information from the assistant. Proactive assistants should make sure to surface information that is relevant to the current context and recent user interactions and messages.
@@
(4) Incorporating user feedback. The proactive assistant should adapt its suggestions and timing as it interacts with the user. The user should be able to accept and reject proactive suggestions, and the decision to accept or reject suggestions should influence future decisions of the proactive assistant. This is especially important as different programmers may want different levels of proactive help from the assistant based on their experience for instance.
(5) Time suggestions based on context. Programmers perform various tasks during their work and are often in a state of flow (Csikszentmihalyi and Larson, 2014). The proactive assistant should consider what the programmer is currently doing and their workflow before it provides a suggestion (Barke et al., 2023). For example, showing suggestions while the user is actively coding can disrupt the programmer’s flow and lead to disruptions. Horvitz (1999) discusses the importance of fallback behavior which can allow for a reduced frequency of suggestions.
@@
In the following section, we introduce an implementation of a proactive assistant for programming by building on the design considerations above.
4. Proactive Assistant Implementation